What is the value of g(x) - f(x) if f(x) = 2x + 4, g(x) = 3x - 2 and x = 12?
3x - 2 - (2x + 4) = 3x - 2 - 2x - 4 = x - 6
If x = 12 , x - 6 = 12 - 6 = 6

Answer by Alan3354(69443)

What is the value of g(x) - f(x) if f(x) = 2x + 4, g(x) = 3x - 2 and x = 12?
3x - 2 - (2x + 4) = 3x - 2 - 2x - 4 = x - 6
If x = 12 , x - 6 = 12 - 6 = 6
--------------------------------
g(12) = 34
f(12) = 28
g(12) - f(12) = 6